### v3

# DiLA: Enhancing LLM Tool Learning with Differential Logic Layer

Abstract
Considering the challenges faced by large language models (LLMs) in logical reasoning and planning, prior efforts have sought to augment LLMs with access to external solvers. While progress has been made on simple reasoning problems, solving classical constraint satisfaction problems, such as the Boolean Satisfiability Problem (SAT) and Graph Coloring Problem (GCP), remains difficult for off-the-shelf solvers due to their intricate expressions and exponential search spaces. In this paper, we propose a novel differential logic layer-aided language modeling (DiLA) approach, where logical constraints are integrated into the forward and backward passes of a network layer, providing another option for LLM tool learning. In DiLA, LLM aims to transform the language description to logic constraints and identify initial solutions of the highest quality, while the differential logic layer focuses on iteratively refining the LLM-prompted solution. Leveraging the logic layer as a bridge, DiLA enhances the logical reasoning ability of LLMs on a range of reasoning problems encoded by Boolean variables, guaranteeing the efficiency and correctness of the solution process. We evaluate the performance of DiLA on three classic constraint satisfaction problems and empirically demonstrate its consistent outperformance against existing prompt-based and solver-aided approaches.
DiLA: Enhancing LLM Tool Learning with Differential Logic Layer
Yu Zhang1  Hui-Ling Zhan2  Zehua Pei1  Yingzhao Lian2  Lihao Yin2 Mingxuan Yuan2Bei Yu1 1The Chinese University of Hong Kong  2Noah’s Ark Lab, Huawei
1 Introduction
Recently, a significant research thrust has been on leveraging large language models (LLMs) for reasoning and planning, with numerous efforts aimed at augmenting their reasoning capabilities. These endeavors include text-based temporal reasoning (Xiong et al., 2024), logic feedback-enhanced alignment methods (Nguyen et al., 2023), and prompt-based methods such as the chain of thoughts (CoT) (Wang et al., 2023) or a simple directive like “Let’s think step by step” (Lightman et al., 2023). Despite the strides made by LLMs in achieving human-like reasoning abilities, they still encounter challenges when confronted with complex plannings (Valmeekam et al., 2022). Occasionally, LLMs exhibit unfaithful reasoning, leading to derived conclusions that do not consistently follow the previously generated reasoning chain in practical applications (Pan et al., 2023).
To this end, recent works have begun to augment LLMs with access to external solvers, i.e., utilize LLMs to first parse natural language logical questions into symbolic representations and subsequently employ external solvers to generate answers based on these representations. To enhance parsing accuracy, LoGiPT (Feng et al., 2023) has been proposed to directly emulate the reasoning processes and mitigate the parsing errors by learning to strict adherence to solver syntax and grammar. It is fine-tuned on a constructed instruction-tuning dataset derived from revealing and refining the invisible reasoning process of deductive solvers. An alternative approach for improving reasoning capabilities involves LOGIC-LM, a new symbolic solver-aided modeling approach (Pan et al., 2023), in which an LLM is used to generate a symbolic problem formulation with in-context learning and offload the actual reasoning task to off-the-shelf symbolic reasoners, such as Z3 solver (De Moura and Bjørner, 2008). This solver-augmented method has demonstrated superiority over CoT by achieving a large margin of $18.4\%$ on average, establishing a new state-of-the-art.
Figure 1: Illustration of CoT (left), solver-aided approach (middle), and our logic layer-aided language modeling approach (right).
Despite their impressive performance on various benchmark tests, LOGIC-LM (Pan et al., 2023) and its solver-augmented successors can not deal with reasoning and planning problems in practical scenarios. Particularly, these real problems often state a set of premises and complex constraints and require a sophisticated search process to find the optimal solution, which is still challenging even for modern solvers. More specifically, even the state-of-the-art symbolic solvers, such as Z3(De Moura and Bjørner, 2008) and Kissat(Biere and Fleury, 2022), encounter significant bottlenecks when addressing formal verification problems such as electronic circuit verification involving tree/cyclic circuit structures(Shi et al., 2023) and job-shop scheduling problems (Li et al., 2022) with complex graph connections. For instance, the Kissat solver requires days or even weeks to process a vanilla circuit verification problem with arithmetic circuit modules including multipliers or multiply-add circuits, yet fails to produce a solution. This highlights the limitations of solver-augmented LLM approaches which solely rely on symbolic solvers to tackle logical reasoning problems in reality.
In this paper, we propose DiLA, a novel tool-learning approach for Large Language Models, designed to enhance their logical reasoning capabilities through the integration of an additional logic layer. Unlike existing methods that either rely on in-context prompting for step-by-step reasoning (see Figure 1(left)) or entirely offload reasoning to external solvers (see Figure 1(middle)), our approach offers a third option: enhancing LLMs’ reasoning ability by incorporating a differential logic layer (see Figure 1(right)). Specifically, DiLA leverages the LLM to parse and comprehend the problem description, generate an initial solution based on its language understanding, and then iteratively refine this solution through forward and backward passes of a network layer—the differential logic layer—which embeds first-order logic constraints into its architecture. In this way, DiLA overcomes the limitations of traditional solvers by directly performing reasoning within the framework of layer-augmented LLMs. Our contributions are summarized as follows:
  • We introduce a novel tool-learning approach for LLMs, DiLA, which synergistically integrates a differential logic layer into LLMs, effectively bridging the gap between natural language understanding and symbolic reasoning capabilities.
  • Leveraging SAT encoding as a bridge, DiLA successfully translates natural language reasoning problems into satisfiability problems, enabling it to tackle a range of reasoning problems, such as SAT and GCP.
  • Compared to other solver-aided LLM approaches, DiLA pioneers a novel strategy that differentiates symbolic problems and iteratively searches for solutions through forward and backward propagation of a network layer, thereby circumventing the limitations of off-the-shelf solvers.
We evaluate the performance of our approach on three constraint satisfaction problems: logical deduction, Boolean satisfiability, and graph coloring. Our analysis yields two key findings: firstly, on simple artificial problem instances, DiLA boosts the inference accuracy of LLMs to 100% and consistently outperforms solver-aided approaches with improved runtime. Secondly, for real-world problems that current solvers struggle with, DiLA showcases robustness and remarkable efficiency in handling these complex test cases, thereby opening up opportunities for further real-world applications.
2 Motivation
Figure 2: Exemplar comparison of solving graph coloring problems by different approaches. Direct prompts by GPT-4 make errors when generating the color assignment step by step; LOGIC-LM based on the Z3 symbolic solver cannot solve parsed constraints and outputs unknown (both errors are highlighted in red). In contrast, DiLA can generate the correct answer by combining the strengths of LLMs and the differential logic layer.
This paper explores the potential of language understanding and logical reasoning capabilities in LLMs. Traditionally, the prevailing approaches have either relied solely on LLMs for step-by-step reasoning (Wang et al., 2023; Zhou et al., 2022) or offloaded reasoning tasks to off-the-shelf solvers (Olausson et al., 2023; Ye X, 2023). However, we propose a third approach, as these two extremes either underutilize or over-rely on LLMs during reasoning. Specifically, DiLA leverages the powerful understanding abilities of LLMs to extract logical formulas and generate possible solutions. By utilizing LLMs, such as GPT-4, as a solution generation engine, we can produce an initial solution based on input semantic constraints, laying the groundwork for further refinement.
The second motivation for this research stems from an analysis of current solver applications in real-world scenarios, characterized by two distinctive features: (1) The large scale of reasoning problems in reality, leading to rapid degradation in the performance of heuristic-based solvers due to the exponential expansion of the search space; (2) The formidable challenge presented by the complex structure of logic formulas, often requiring weeks or even months to resolve using current solvers. Both characteristics significantly limit the effectiveness of a solver-augmented LLM in the domain of logic reasoning.
Consequently, our objective is to identify a synergistic approach that combines the strengths of LLMs and a differential logic layer, thereby circumventing the limitations of traditional solvers. This approach leverages the LLM’s capacity to comprehend logical formulas while concurrently utilizing the logic layer’s refinement abilities to achieve accurate solutions. Figure 2 illustrates a comparison between CoT, LOGIC-LM, and our proposed DiLA. The LLM alone may introduce logical flaws during step-by-step inference, such as assigning the same color to vertex 1 and vertex 4 despite their edge connection, and LOGIC-LM may struggle with complex reasoning problems due to its backbone solver’s limitations. In contrast, DiLA produces accurate answers through the collaboration of the LLM and the differential logic layer. Specifically, DiLA first uses the LLM to parse a natural language input into logic constraints and generate an initial solution based on its semantic understanding, then employs the differential logic layer to refine this initial solution.
3 Differenial Logic Layer-Aided Language Models
3.1 Overview
Figure 3: The overall flow of DiLA. Starting from a problem in natural language, in Step 1, the LLM parses it into a SAT problem specification which consists of a set of logical formulas. Then in Step 2, the backbone LLM tries to output an initial solution based on its language understanding. In Step 3, the relaxed variables and extracted constraints are offloaded to a differential logic layer, checking the constraint satisfiability, and updating the solution until until all constraints are met. Here, the solid arrow indicates the forward pass, while the dashed arrow represents the backward pass.
In this section, we present DiLA, which augments LLM with the ability of logical reasoning by incorporating a differential logic layer. More specifically, DiLA addresses the challenge of using LLMs to tackle canonical reasoning tasks expressed in natural language. These tasks typically involve presenting a set of premises and constraints, prompting questions that necessitate intricate deductive reasoning over the provided inputs, which remains a formidable challenge even for contemporary LLMs (Valmeekam et al., 2022).
The general procedure for solving natural language reasoning tasks with DiLA can be conceptualized in three distinct steps: parsing, initialization, and refinement. Given a natural language input that describes both the propositional constraints $\phi$ and the question $Q$, we first parse this input into a SAT specification using LLMs, thereby obtaining a formal description of the constraints and variables. Next, we leverage the LLM’s natural language understanding to generate an initial variable assignment. Since this initial solution may only partially satisfy the propositional constraints, DiLA iteratively refines it through a differential logic layer that encodes all of the logical formulas, ultimately yielding a more accurate solution. Figure 3 illustrates the overall flow of our proposed DiLA.
3.2 Problem Formulator and Initialization
Intuitively, LLMs may struggle with directly solving complex reasoning problems. However, they have demonstrated a notable ability to comprehend textual inputs and translate them into formal programs, such as mathematical equations (He-Yueya et al., 2023) or satisfiability modulos (Ye X, 2023). Notably, the SAT problem can serve as a versatile intermediate step for solving a broad range of constraint satisfaction problems, provided they can be expressed using Boolean variables. Specifically, problem instances from NP-complete domains, such as Graph Coloring and Set Cover, can be seamlessly encoded into SAT problem specifications, thereby allowing for efficient solutions via SAT algorithms (Stechly et al., 2024; Liu et al., 2023). Therefore, we harness SAT encoding as a general bridge to tackle these constraint satisfaction problems in practical settings.
Specifically, given a problem description in natural language, DiLA prompts an LLM with detailed instructions to generate the SAT specification, which includes a set of premises and constraints. Typically, the SAT specification here involves conjunctive normal forms (CNFs), denoted as $\phi(v_{1},\dots,v_{n})$, which is a conjunction of clauses (constraints) $C$. Formally, rules are written in the conjunctive form of clauses $C_{1}\wedge C_{2}\cdots\wedge C_{m}$, which each $C_{j}$ is a constraint. The total rule is satisfied if and only if all of the clauses $C_{1},C_{2},\dots,C_{m}$ are simultaneously True. Each clause represents a disjunction of literals, where a literal is either a propositional variable $v_{i}$ or its complement $\neg v_{i}$, for example, $v_{1}\vee\neg v_{2}$. In other words, if at least one literal in a clause is True, this clause would also be True. Variables can be assigned logic values, either $1$ or $-1$, representing True or False, respectively  [^1] In other works, they may claim the logic value of each literal is 00 or $1$. It should be noted that the two claims are equal under simple mathematical transformations.. See Appendix A for more details on the SAT specification.
Aside from problem formulation, leveraging an LLM can be a valuable strategy to generate an initial solution. We observe that, after generating the SAT specification, we can prompt LLMs, like GPT-4 (Achiam et al., 2023) or Llama-3 (Meta, 2024), to produce a potential solution or a set of possible solutions. This can be achieved by framing the problem as a natural language query, such as “What is the logical solution based on the premises?” or “Can you provide a possible answer from these given constraints?” (see Appendix B for more details). The LLM’s response can then serve as a starting point for further refinement and validation, allowing us to build upon its output and iteratively improve the solution through logical analysis and reasoning. By harnessing the LLM’s ability to understand semantic constraints, we can tap into its potential to facilitate the initial solution-finding process and accelerate solving progress towards a well-reasoned answer.
3.3 From SAT to Differential MaxSAT Formulations
In fact, SAT is very efficient at expressing constraint satisfaction problems in that many of the standard NP-complete problems, like graph coloring, can be straightforwardly formulated and efficiently solved as SAT problems (Buss and Nordström, 2021; Liu et al., 2023). Traditionally, a SAT solver, such as Z3 and Kissat, is leveraged to determine a satisfying assignment for the given constraint formula $\phi$. Contemporary SAT solvers are founded on the Conflict-Driven-Clause-Learning (CDCL) algorithm, which excels in its ability to learn from conflicts and use that conflict knowledge to prune the branch-and-bound search space more effectively. However, existing CDCL-based SAT solvers still suffer from exponential searching space and are unable to correct errors through a learning-from- mistakes system, resulting in an infinite loop in solving complex SAT problems (Shi et al., 2023).
In this study, when an NL reasoning problem is fed into DiLA, our goal is to determine its solution that can satisfy all logical constraints. To achieve this, we might incorporate a solver as an additional logic reasoner for the LLM, as in SATLM (Ye X, 2023) or LOGIC-LM (Pan et al., 2023). However, the exponential search complexity inherent in heuristic-based SAT solvers poses a significant challenge, limiting their effectiveness when dealing with complex real-world problems. Therefore, a key issue is how to design an efficient solver surrogate that can both be seamlessly integrated into LLMs and efficiently address logical reasoning problems.
The maximum satisfiability (MaxSAT) problem serves as the optimization counterpart to the SAT problem, aiming to maximize the number of satisfied clauses. Indeed, if a solution to the MaxSAT problem can satisfy all the clauses, the variable assignment can be used to constitute a valid solution for the original SAT problem. In the case of the SAT problem, each CNF is associated with a set of clauses (constraints), and each clause is defined on a subset of variables, signifying the variables’ simultaneous legal assignments. Formally, each CNF $\phi(v_{1},\dots,v_{n})$ comprises $n$ binary variables, with each $v_{i}\in\{1,-1\}$ ($i\in{1,2,\dots,n}$) representing a boolean variable. Let’s introduce the coefficient $\boldsymbol{c}_{j}\in\{-1,0,1\}^{m}$, where $c_{ij}$ denotes the sign of $v_{i}$ in clause $j\in{1,2,\dots,m}$. Consequently, we can establish a clause matrix $\boldsymbol{C}\in\{1,-1,0\}^{m\times n}$, where each element $c_{ij}$ in $\boldsymbol{C}$ signifies the sign of variable $\tilde{v}_{i}$ in clause $j$. Therefore, each SAT instance can be translated into a corresponding MaxSAT problem, wherein $\bigvee$ represents the logical “or” symbol,
$\underset{\tilde{\boldsymbol{v}}\in\{-1,1\}^{n}}{\max}\sum_{j=1}^{m}\bigvee_{i=1}^{n}\textbf{1}\{c_{ij}\tilde{v}_{i}>0\}.$ (1)
We further formulate Equation 1 in its minimization, or unsatisfiability, equilibrium as
$\underset{\tilde{\boldsymbol{v}}\in\{-1,1\}^{n}}{\min}\sum_{j=1}^{m}\bigwedge_{i=1}^{n}\textbf{1}\{c_{ij}\tilde{v}_{i}<0\},$ (2)
where $\bigwedge$ is the logical “and” symbol. Indeed, the objective value in Equation 2 is 0 if and only if a satisfiable solution can be found. Our goal is to establish a continuous upper bound, referred to as the “loss”, for each clause to quantify its level of unsatisfiability. In essence, the loss takes an upper bound if the clause is unsatisfied, and by minimizing this loss, we can strive to push it closer to satisfaction. To make a purely quadratic loss function as in  (Wang and Kolter, 2019), we introduce $v_{0}=1$ and $s_{0j}=-1$ in Equation 2. Therefore, the minimization problem in Equation 2 can be solved by transforming into a quadratic loss function as:
$\mathcal{L}_{j}=\frac{(\sum_{i=0}^{n}c_{ij}\tilde{v}_{i})^{2}-(m_{j}-1)^{2}}{4m_{j}},\quad\mathcal{L}=\sum_{j=1}^{m}\mathcal{L}_{j},$ (3)
where $\mathcal{L}_{j}$ is the loss value of $j$-th clause, $\mathcal{L}$ is the loss value of all clauses, and $m_{j}$ is the number of literals in clause $j$, e.g., 3 for the Max3SAT problem. The loss function in Equation 3 is essentially a quadratic loss that takes the upper bound when no literal in clause $j$ is satisfied. In other words, it captures the extent of unsatisfiability for a given clause by penalizing solutions that fail to satisfy any of its literals. Specifically, for any value of $m_{j}$, it can be easily verified that this quantity is equal to +1 if no literal is satisfied, and 0 or less if at least one literal is True.
Now, the MaxSAT solving is equivalent to finding an assignment vector $\tilde{\boldsymbol{v}}\in\{-1,1\}^{n}$ that minimizes loss in Equation 3. By relaxing each discrete variable $\tilde{\boldsymbol{v}}_{i}$ to a continuous variable $\boldsymbol{v}_{i}\in\mathbb{R}$, the quadratic loss function becomes
$\mathcal{L}_{j}=\frac{\|\boldsymbol{V}c_{j}\|^{2}-(m_{j}-1)^{2}}{4m_{j}},$ (4)
which is essentially a convex minimization problem. Therefore, leveraging gradient descent to solve this minimization problem, the gradient computation involves differentiating the loss function in Equation 4 with respect to $v_{i}$. Define this gradient as $\boldsymbol{g_{i}}$, we have
$\boldsymbol{g}_{i}=\boldsymbol{V}\boldsymbol{S}^{\top}\boldsymbol{s}_{i}-\|\boldsymbol{s}_{i}\|^{2}\boldsymbol{v}_{i},$ (5)
where $\boldsymbol{S}=[\boldsymbol{c}_{0},\boldsymbol{c}_{1},\dots,\boldsymbol{c}_{n}]\text{diag}(1/\sqrt{4m_{j}})\in\mathbb{R}^{m\times(n+1)}$ and $s_{i}\in\mathbb{R}^{(n+1)}$ is the $i$-th vector in $\boldsymbol{S}$.
Figure 4: Examplar illustration of VCG.
3.4 Differential Logic Layer
We envision the logic layer being used primarily at the top of LLMs, embedding logical formulas $\phi$ produced by the backbone LLM, taking LLM-initialized variable assignments as inputs, and producing outputs that are consistent with $\phi$. Specifically, we draw an analogy between the fully connected (FC) layer and the variable-clause graph (VCG), as depicted in Figure 4, where the weight for positive literals in the clause is 1 (solid arrow) and negative literals in the clause is -1 (dashed arrow). Specifically, we map each variable to an input neuron in an FC layer, each clause to an output neuron, and the coefficients to the weights in the linear transformation with a zero bias vector. While this analogy holds, there are two crucial differences between traditional FC layers and our proposed logic layer. Firstly, the differential logic layer has no unknown parameters, whereas FC layers require data-driven training to learn their weights. Secondly, each clause in the logic layer is only partially connected to variables, in contrast to fully connected layers, where all input neurons are connected to all output neurons.
Supposing that the current solution of MaxSAT in Equation 1 is given as $\phi^{\prime}=\phi(v_{1},v_{2},\dots,v_{n})$, and it is easy to check whether this solution, i.e., variable assignment, can satisfy the original SAT problem $\phi$, which is essentially the forward pass of our logic layer. A trivial case is that if all the clauses (i.e., constraints) are satisfied by the assignments $\mathcal{C}(v_{1},v_{2},\dots,v_{n})$, then these assignments constitute a valid solution for the original SAT problem $\phi$. In most scenarios where $\phi^{\prime}$ satisfies only a subset of the clauses, we define the unsatisfied clauses as $\bar{\phi^{\prime}}$, a subset of all clauses, and denote the indices of the variables involved in $\bar{\phi^{\prime}}$ as $\bar{I}$. Intuitively, the variables in $\bar{I}$ are likely to be the source of conflicts, so we select the variable with the largest absolute gradient from the candidate set $\bar{I}$ and update its value during the backward pass, thereby pushing $\phi^{\prime}$ towards satisfying more constraints. We now elaborate on the forward and backward pass of our proposed logic layer, providing a detailed explanation of its operation.
Forward Pass. The forward pass algorithm is outlined in Algorithm 1. In the forward pass, the inputs consist of relaxed variable assignment at $k$-th iteration. Subsequently, the layer transforms these inputs by extracting the sign of the variables, thereby casting them to Boolean values. The layer then assesses the satisfiability of $\phi^{\prime}$ (line2). If the current variable assignment satisfies $\phi^{\prime}$, the logic layer outputs $y^{k}$ as True, indicating that $\phi$ is satisfied and a feasible solution for the given CNF has been identified. Conversely, if $\phi$ cannot be satisfied, the logic layer outputs $y^{k}$ as False, prompting the initiation of the backward pass to update the variable assignment.
Algorithm 1 The forward pass of logic layer
0:  Solution $\boldsymbol{v}^{k}\in\mathbb{R}^{n}$ at $k$-th epoch.
0:  $y^{k}$, final solution $\boldsymbol{v}^{*}$.
1:  $\tilde{\boldsymbol{v}}^{k}\leftarrow[\boldsymbol{v}^{k}_{i}>0:i=1,\dots,n]$;
2:  $\phi^{\prime}\leftarrow\phi(\tilde{v}^{k}_{1},\dots,\tilde{v}^{k}_{n})$;
3:  if $\phi^{\prime}$ is satisfiable then
4:     $y^{k}\leftarrow$ True;
5:     $\boldsymbol{v}^{*}\leftarrow\tilde{\boldsymbol{v}}^{k}$ ;
6:  else
7:     $y^{k}\leftarrow$ False;
8:  end if
Backward Pass. The backward pass is responsible for computing the gradients of the layer inputs and derives updates to variables that steer towards satisfying the constraints $\phi$. A crucial aspect of the backward pass is identifying the input variables that contribute most to the unsatisfiability of the constraint formulas. It is well-established that variables in $\bar{I}$ form the unsatisfiable subset and are, therefore, more likely to be sources of conflict. Conversely, variables not present in $\bar{I}$ can have their gradients set to zero, as their absence in the conflict clauses provides no evidence regarding the correctness or incorrectness of their values. Inspired by the stochastic local search (SLS) algorithm, commonly used in constraint satisfaction problems (Chu et al., 2023), we select the variable with the largest absolute gradient from the candidate set $\bar{I}$ and update its value at each iteration. However, our logic layer diverges from SLS in that it employs a “differential” variable selection mechanism during backpropagation, whereas SLS relies on meta-heuristics.
Algorithm 2 illustrates our backward pass. The backward pass begins by initializing the gradient to zero for all variables (line 1). If $y^{k}$ is false, indicating the presence of unsatisfied clauses, we obtain the set of variables $\bar{I}$ that are present in the falsified clauses $\bar{\phi^{\prime}}$. Once we have obtained the candidate set $\bar{I}$ (line 5), we proceed to select the best variable from this set based on its gradient. Specifically, we compute the gradient as in Equation 5 for each variable in the candidate set (line 7). Then, the logic layer selects a variable and updates its value based on two situations: (1) If there exists a variable with a non-zero gradient (i.e., $g_{i}\neq 0$), the variable with the largest absolute $g_{i}$ would be selected (line 10); (2) If there is no variable satisfying the above condition, indicating that the search is stuck in a local optimum, we randomly select a variable from a falsified clause (line 12) and artificially assigns a gradient that would change its sign after gradient descent (line 13). This criterion guides the updates towards satisfying more clauses at each iteration, as selecting the variable with the largest absolute gradient pushes the loss quantity to decrease in the steepest direction. The forward and backward propagation iterates until the logic layer finds a satisfying assignment or exceeds the maximum time limit.
Algorithm 2 The backward pass of logic layer
0:  $\boldsymbol{v}^{k}\in\mathbb{R}^{n}$ from forward pass, $y^{k}$ from forward pass, learning rate $\lambda$.
0:  Gradient $\boldsymbol{g}^{k}$ of $\boldsymbol{v}^{k}$, updated assignment $\boldsymbol{v}^{k+1}$.
1:  $\boldsymbol{g}^{k}\leftarrow\emph{0}$;
2:  if $y^{k}$ is False then
3:     $\tilde{\boldsymbol{v}}^{k}\leftarrow[\boldsymbol{v}^{k}_{i}>0:i=1,\dots,n]$;
4:     $\phi^{\prime}\leftarrow\phi(\tilde{v}^{k}_{1},\dots,\tilde{v}^{k}_{n})$;
5:     $\bar{I}\leftarrow\{i\in[1,n]|v_{i}\in\bar{\phi^{\prime}}\}$;
6:     for $i\in\bar{I}$ do
7:        $g_{i}\leftarrow\partial_{\boldsymbol{v}^{k}_{i}}\mathcal{L}$;
8:     end for
9:     if $\exists\;g_{i}\neq 0$ then
10:        $g^{k}_{i}\leftarrow\arg\max_{i\in\bar{I}}\|g_{i}\|$
11:     else
12:        $v_{i}:=$ a random variable in a falsified clause;
13:        $g^{k}_{i}:=\text{sign}(v_{i})$;
14:     end if
15:  end if
16:  Update $\boldsymbol{v}^{k+1}\leftarrow\boldsymbol{v}^{k}-\lambda\boldsymbol{g}^{k}$;
4 Experiments
Table 1: Accuracy and runtime (s) of CoT, including GPT-4 and Llama-3, LOGIC-LM, and DiLA on LogicalDeduction, simple Boolean Satisfiability and 3-Coloring datasets.
Problem Number of Variables GPT-4 Llama-3 LOGIC-LM DiLA
Acc.(%) Acc.(%) Acc.(%) Time(s) Acc.(%) Time(s)
Logical
Deduction
3,5,7 76 81 100 0.01 100 0.01
SAT 20 12 17 100 0.05 100 0.01
50 5 0 100 0.12 100 0.01
100 0 0 100 0.17 100 0.03
200 0 0 100 0.99 100 0.04
250 0 0 100 5.24 100 0.08
Graph Coloring 10 20 22 100 0.05 100 0.01
50 7 10 100 0.19 100 0.02
100 0 0 100 0.58 100 0.04
150 0 0 100 2.64 100 0.15
200 0 0 100 5.70 100 0.24
4.1 Setup
Tasks: We conduct experiments on three fundamental reasoning tasks: the logical deduction problem, Boolean Satisfiability, and Graph Coloring problems. The logic deduction problems are mostly about deducing the order of a sequence of objects from a minimal set of conditions. Here we utilize LogicalDeduction dataset from the BigBench (Srivastava et al., 2022) collaborative benchmark. For Boolean Satisfiability problems, we utilize open-source benchmark instances  [^2] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html with 20 to 250 variables, focusing on finding variable assignments that satisfy all constraints. Notably, we do not consider the unsatisfiable cases as finding unsatisfied certificates is a distinct problem. For Graph Coloring problems, we randomly generated 100 3-coloring instances with vertices counts ranging from 10 to 200, following the approach in  (Stechly et al., 2024) (see Appendix C for details), aiming to color graph vertices so that no two adjacent vertices share the same color. Furthermore, we also incorporate several complex SAT cases in SAT Competition 2023333https://satcompetition.github.io/2023/ and open-source graph coloring problems with large number of edges [^4] https://www.cs.ubc.ca/~hoos/SATLIB/Benchmarks/SAT/ to illustrate the robustness of our proposed DiLA. These instances originate from practical industrial problems and frequently pose challenges for modern SAT solvers. Notably, current solvers often enter into an infinite loop when handling these test cases due to their complex structure, requiring a prohibitively long time to solve.
Baselines: We conducted a comparative analysis between DiLA and three baselines, including the prompt-based methods with GPT-4 (Achiam et al., 2023) and Llama-3 (Llama-3-70B-Instruct) (Meta, 2024), and the solver-augmented method, LOGIC-LM (Pan et al., 2023). In particular, LOGIC-LM (Pan et al., 2023) employs GPT-4 to parse problem specifications and offloads the logical reasoning task to the symbolic solvers, including CSP Solver [^5] https://github.com/python-constraint/ python-constraint and Z3 solver (De Moura and Bjørner, 2008), which serves as the state-of-the-art logic reasoning methods.
Setup: We implement a prototype of our proposed DiLA using Pytorch (Paszke et al., 2019), leveraging GPT-4 as the backbone LLM model. Notably, the logic layer within DiLA has no training parameters and can adapt to various problem types expressed in Boolean variables. We employ the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $2\times 10^{-1}$ in DiLA, facilitating the effective updating of selected variables. Furthermore, we use a temperature of 0 for LLMs, consistent with the LOGIC-LM approach. We set a time limit of 10,000 seconds for both the solvers and the logic layer.
4.2 Main Results
We report the accuracy of DiLA and the three baselines in Table 1. Accuracy is evaluated based on whether the LLM can output a correct answer that satisfies all constraints. We evaluate LLMs over 100 instances in each domain. In addition to accuracy, we also report the solving runtime for LOGIC-LM and DiLA, both of which leverage GPT-4 as a backbone LLM to perform language understanding. Analysis of Table 1 reveals that both solver-augmented LOGIC-LM and logic layer-aided DiLA achieve 100% accuracy on these simple benchmarks, owing to the precise logic parsing and accurate solving. In contrast, standalone LLMs, including GPT-4 and Llama-3, cannot handle reasoning problems with over 100 variables with CoT prompting. More importantly, for all test instances, DiLA exhibits faster performance than LOGIC-LM in solving process, especially for relatively large cases with over 200 variables. The runtime speedup can be up to 65.5$\times$ when dealing with SAT problems with 250 variables and 23.8$\times$ when tackling 3-coloring problems with 200 nodes. Overall, both the solver-augmented LLM and our proposed DiLA can successfully address these simple artificial test cases, with the runtime speedup highlighting the efficiency of DiLA in logic reasoning.
Table 2: Runtime (s) of Z3, Kissat, and our DiLA on solving hard reasoning problems.
Problem Test case #Variables #Clauses Z3 Kissat DiLA
SAT rbsat 1150 84314 >10,000 >10,000 98.76
sgen3 260 884 >10,000 >10,000 11.83
Schur 756 28445 >10,000 >10,000 23.81
SCPC 900 41714 >10,000 >10,000 20.23
Graph Coloring g125.17 125*17 68397 >10,000 >10,000 29.57
g125.18 125*18 72413 >10,000 >10,000 31.14
g250.15 250*15 237715 >10,000 >10,000 29.68
g250.29 250*29 461872 >10,000 >10,000 30.07
4.3 Robustness of Reasoning
Compared with SATLM (Ye X, 2023) and LOGIC-LM (Pan et al., 2023), which rely on an off-the-shelf solver for reasoning, DiLA sidesteps the limitations of the solvers themselves, such as high computational costs when dealing with complex formulas. In practice, there are intricate cases where even state-of-the-art SAT solvers struggle, sometimes taking weeks to solve. These challenging instances often involve Boolean satisfiability formulas with complex structures and graph coloring problems with considerable edge connections. For our evaluation, we leverage the state-of-the-art SMT solver, Z3 (De Moura and Bjørner, 2008), and the widely-used SAT solver, Kissat (Biere and Fleury, 2022), as baselines, and test on some hard test cases derived from real-world problems.
The experimental results are shown in Table 2, revealing several key observations. First, DiLA successfully solves these challenging cases within a reasonable runtime, whereas both Z3 and Kissat fail to produce valid results within the time limit. As indicated in Table 2, these difficult constraint satisfaction instances typically exhibit large clause-to-variable (CV) ratios (m/n). For example, the test case “rbsat” consists of only 1150 variables, but the total number of clauses amounts to 84,314, resulting in an extremely high CV ratio of 73.32. This suggests that there is likely only one viable solution for these problems. If a traditional solver’s initial search path deviates significantly from the correct path, the CDCL framework’s inability to rectify errors through a learning-from-mistakes system can lead to an endless loop. In contrast, our proposed DiLA, initialized by LLMs and guided by differentiation of the loss function, enables us to first reach a partially satisfied solution and then progressively update it using an efficient searching strategy, rather than completely failing if stuck.
Second, the initial solution generated by backbone LLMs can serve as an excellent starting point. For example, in the SAT case “SCPC”, after analyzing all constraints, the backbone LLM provides an initial solution with an unusual all-false variable assignment. We discovered that over 99% of the final satisfying variable assignments should indeed be set to 0, indicating that a significant proportion of variables require no further updates. Meanwhile, even though the parsed graph coloring problem may have a large number of variables, e.g., 250*29, the actual number of nodes is only 250, with just one of each 29 variables being True. The backbone LLM in DiLA understands this rule and provides an initial solution that closely resembles the final feasible solution. In contrast, traditional solvers rely on random initialization and need to explore all possible assignments for each variable before reaching the final solution, resulting in exponential search spaces and reduced efficiency. In general, the evaluation results on real-world cases demonstrate the high efficiency of our proposed DiLA.
5 Related Works
Prior approaches to NL-based reasoning with LLMs can be broadly categorized into two groups. One is in-context learning approaches that design special prompts to elicit LLMs’ step-by-step reasoning capabilities. Typical methods include chain-of-thought prompting (Wang et al., 2023) that generates a sequence of reasoning steps before the final answer and the least-to-most prompting (Zhou et al., 2022) that breaks the problem down into simpler components that can be solved sequentially. Both the above approaches perform reasoning directly over natural language (NL), providing greater flexibility than symbolic-based reasoning. However, the intrinsic complexity and ambiguity of NL also bring undesired issues such as unfaithful reasoning and hallucinations. The other is tool-augmented approaches that only require LLMs to parse the problem specification out of the language description accurately and leverage off-the-shelf automated tools to derive the final answer, as exemplified by SATLM (Ye X, 2023) and LOGIC-LM (Pan et al., 2023). The tool-augmented approaches guarantee the correctness of the answer with respect to the parsed specification and avoid planning errors in the solving process. However, the performance of such tool-augmented methods highly relies on external tools and can not deal with many real problems due to the deficiency of backbone solvers.
6 Conclusion
In this work, we introduce a pioneering method named differential logic layer-aided language modeling (DiLA). Starting with an NL reasoning problem, DiLA first uses an LLM to cast it into a SAT problem and generate a possible solution based on its language understanding, and then progressively refines this solution within a logic layer. In this way, we harness the potential of the language-understanding ability of LLMs and sidestep the limitations of off-the-shelf solvers. Extensive experiments on two reasoning tasks demonstrate the superior efficiency of our approach over state-of-the-art solver-augmented LLMs. Notably, DiLA demonstrates its robustness in dealing with challenging problem instances that SOTA solvers struggle with. The results suggest that DiLA achieves new state-of-the-art performance in symbolic logical reasoning tasks, paving the way for more applications of LLMs in practical reasoning settings.
7 Limitation
We identify two main limitations of DiLA. First, DiLA relies on translating reasoning problems into logical formats that can be encoded by boolean variables and first-order logic. As a consequence, the model’s applicability is inherently bounded by the expressiveness of the underlying theory, as not all problems can be easily encoded in first-order formulas with boolean vectors. Nevertheless, the wide range of tasks that we can instantiate our DiLA framework on shows its general applicability.
Second, DiLA relies on backbone LLMs to translate natural language problems into symbolic representations. Although this approach has shown promise, it may struggle with logical representations featuring complex grammar structures, such as probabilistic circuit logic. This limitation stems from the challenge of communicating intricate grammatical rules to the language model through a limited number of demonstrations within a constrained context size. To address this, future research could investigate the development of specialized modules to improve the mapping between natural language and symbolic language, for instance, by fine-tuning LLMs with symbolic formulas for specific reasoning tasks.
Appendix A Details of the SAT Specification
To better leverage the parametric knowledge that LLMs have acquired from pretraining on vast amounts of language data, our approach uses a SAT specification as a bridge to encode a range of semantic constraints. Specifically, it translates general reasoning problems into formal constraints with Boolean variables. Below, we provide a prompt example of how to convert a classical graph 3-coloring problem into its SAT specification in Table 3.
Example prompt of logic formula generation
Question: Color the following graph with 3 colors such that no two vertices share the same color. Vertex 0 is connected to vertex 4. Vertex 0 is connected to vertex 5. Vertex 1 is connected to vertex 4. $\cdots$ Vertex 8 is connected to vertex 5. Vertex 9 is connected to vertex 4. Vertex 9 is connected to vertex 5. There are a total of 10 vertices. What is the SAT formulation for this graph coloring problem?
LLM response: The SAT (Satisfiability) formulation for the graph coloring problem involves encoding the problem as a Boolean satisfiability problem. Here’s the SAT formulation for the graph coloring problem:
Variables:
For each vertex v and each color c (Red, Blue, Green), we introduce a Boolean variable x_vc that represents whether vertex v is assigned color c.In this case, we have 10 vertices and 3 colors, so we introduce 30 Boolean variables: x_0_Red, x_0_Blue, x_0_Green,…, x_9_Red, x_9_Blue, x_9_Green.
Clauses:
We need to encode the following constraints: 1. Each vertex must have exactly one color: For each vertex v, we add a clause that ensures exactly one color is assigned:
$(x\_v\_Red\vee x\_v\_Blue\vee x\_v\_Green)\wedge\neg(x\_v\_Red\vee x\_v\_Blue)\wedge\neg(x\_v\_Red\vee x\_v\_Green)$
$\wedge\neg(x\_v\_Blue\vee x\_v\_Green)$
This clause ensures that vertex v has at least one color (first part) and at most one color (second to fourth parts).
2. Adjacent vertices must have different colors: For each edge (u, v) in the graph, we add a clause that ensures the vertices have different colors:
$\neg(x\_u\_Red\vee x\_v\_Red)\wedge\neg(x\_u\_Blue\vee x\_v\_Blue)\wedge\neg(x\_u\_Green\vee x\_v\_Green)$
This clause ensures that vertices u and v do not have the same color.
Complete SAT formulation:
The complete SAT formulation consists of the conjunction of all the clauses:
$(x\_0\_Red\vee x\_0\_Blue\vee x\_0\_Green)\wedge\neg(x\_0\_Red\vee x\_0\_Blue)\wedge\neg(x\_0\_Red\vee x\_0\_Green)\wedge\neg(x\_0\_Blue\vee x\_0\_Green)\wedge$
$\cdots$
$(x\_9\_Red\vee x\_9\_Blue\vee x\_9\_Green)\wedge\neg(x\_9\_Red\vee x\_9\_Blue)\wedge\neg(x\_9\_Red\vee x\_9\_Green)\wedge\neg(x\_9\_Blue\vee x\_9\_Green)\wedge$
$\neg(x\_0\_Red\vee x\_4\_Red)\wedge\neg(x\_0\_Blue\vee x\_4\_Blue)\wedge\neg(x\_0\_Green\vee x\_4\_Green)\wedge$
$\cdots$
$\neg(x\_9\_Red\vee x\_5\_Red)\wedge\neg(x\_9\_Blue\vee x\_5\_Blue)\wedge\neg(x\_9\_Green\vee x\_5\_Green)$
Table 3: Prompt example of converting a classical graph 3-coloring problem into its SAT specification.
We can then leverage the transformed logic formula to construct the proposed logic layer. Specifically, for large graphs with many nodes and edges, we ask the LLM to generate Python code to help transform the natural language description into a SAT specification, ensuring the process is both quick and accurate.
Appendix B Details of the Initialization
After parsing logic formulas from the problem description, we can directly ask LLM to generate a possible solution according to its language understanding. The possible question prompt can be “What is the logical solution based on the premises?” or “Can you provide a valid answer from these given constraints?”. Here we show a prompt example for a simple boolean satisfiability problem in Table 4.
Example prompt of initial solution generation
Question: Consider three Boolean variables: A, B, and C. The following constraints must be satisfied: 1. At least one of A or B is true; 2. At least one of B or C is true; 3. A and C cannot both be true; 4. B cannot be true if A is true. Can you analyze this Boolean satisfiability problem and give a valid answer?
LLM response: Let’s break down the constraints:
1. At least one of A or B is true: $A\vee B$.
2. At least one of B or C is true: $B\vee C$.
3. A and C cannot both be true: $\neg(A\wedge C)=\neg A\vee\neg C$.
4. B cannot be true if A is true: $\neg(A\wedge B)=\neg A\vee\neg B$.
Therefore, the overall constraint formula would be $(A\vee B)\wedge(B\vee C)\wedge(\neg A\vee\neg C)\wedge(\neg A\vee\neg B).$
Now, let’s analyze these constraints:
From constraint 4, we know that if A is true, B must be false. This means that if A is true, the only way to satisfy constraint 1 is to have B false and C true (since at least one of B or C must be true).
From constraint 3, we know that A and C cannot both be true. This means that if A is true, C must be false.
So, we have two possible scenarios:
Scenario 1: A is true, B is false, and C is false.
Scenario 2: A is false, B is true, and C is true.
Let’s check if these scenarios satisfy all the constraints:
Scenario 1:
1. Constraint 1: $A\vee B$ is true (A is true).
2. Constraint 2: $B\vee C$ is false (both B and C are false), but this is okay because we can still satisfy it by making C true.
3. Constraint 3: $\neg A\vee\neg C$ is true (A is true, C is false).
4. Constraint 4: $\neg A\vee\neg B$ is true (A is true, B is false).
Scenario 2:
1. Constraint 1: $A\vee B$ is true (B is true).
2. Constraint 2: $B\vee C$ is true (B is true).
3. Constraint 3: $\neg A\vee\neg C$ is true (A is false, C is true).
4. Constraint 4: $\neg A\vee\neg B$ is true (A is false, B is true).
Both scenarios satisfy all the constraints. Therefore, the Boolean satisfiability problem has two valid solutions:
1. A = true, B = false, C = false
2. A = false, B = true, C = true
These are the two possible assignments of values to A, B, and C that satisfy all the constraints.
Table 4: Example of generating an initial solution for a boolean satisfiability problem.
From this prompt response by the LLM, we have two key observations. First, the LLM can successfully understand semantic constraints, transform them into logical formulas, and perform logical reasoning. The possible solution, i.e., variable assignment, provided by the backbone LLM is reasonable and can satisfy some of the given constraints based on its language understanding. However, LLMs still make some mistakes when performing logical reasoning. For instance, in Scenario 1, the solution provided does not satisfy all constraints, despite the LLM’s assertion that it does (as indicated by the words in red in the response). Therefore, it is necessary to design a powerful tool to check the validity of the given solution and further refine it until all constraints are satisfied.
Appendix C Dataset Generation
Graph coloring benchmarks belong to the class of NP-complete problems with which exact heuristic-based solvers struggle. To test the effectiveness of our proposed DiLA, we built our dataset using GrinPy [^6] https://pypi.org/project/grinpy/ for common graph operations. Once a successful candidate is found, it is compiled into the standard DIMACS format and appended with a comment containing its precalculated chromatic number. Specifically, before inputting to LLMs, we transform the graph connections in DIMACS format (e.g., 0 3) into natural language statements (e.g., Vertex 0 is connected to vertex 3.). For the experiments in Section 4.2, we generated 100 graph coloring instances with node counts ranging from 10 to 200 and edge counts from 30 to 480. Notably, all these artificial graphs can be successfully colored with 3 colors, providing an ideal and easily solvable setting. An example of one of the graphs with 10 nodes and 30 edges as well as its 3-coloring solution is shown in Figure 5.
(a) Graph instance with 10 nodes
(b) Graph instance after coloring
Figure 5: 3-coloring graph example.
We also list the semantic edge connections for this 3-coloring problem in Table 5. After explaining the edge connections given in DIMACS format, LLMs can directly interpret the pure numbers, such as “0 3”, as edge connections, eliminating the need to generate long language sequences for large-scale problems.
Example description for edge connection
Vertex 0 is connected to vertex 3. Vertex 0 is connected to vertex 4.
Vertex 0 is connected to vertex 5. Vertex 1 is connected to vertex 3.
Vertex 1 is connected to vertex 8. Vertex 1 is connected to vertex 9.
Vertex 2 is connected to vertex 3. Vertex 3 is connected to vertex 0.
Vertex 3 is connected to vertex 1. Vertex 3 is connected to vertex 2.
Vertex 3 is connected to vertex 4. Vertex 3 is connected to vertex 5.
Vertex 3 is connected to vertex 6. Vertex 3 is connected to vertex 7.
Vertex 3 is connected to vertex 8. Vertex 3 is connected to vertex 9.
Vertex 4 is connected to vertex 0. Vertex 4 is connected to vertex 3.
Vertex 4 is connected to vertex 7. Vertex 5 is connected to vertex 0.
Vertex 5 is connected to vertex 3. Vertex 6 is connected to vertex 3.
Vertex 7 is connected to vertex 3. Vertex 7 is connected to vertex 4.
Vertex 7 is connected to vertex 8. Vertex 8 is connected to vertex 1.
Vertex 8 is connected to vertex 3. Vertex 8 is connected to vertex 7.
Vertex 9 is connected to vertex 1. Vertex 9 is connected to vertex 3.
Table 5: Edge connections in language for a 3-coloring problem.